# CoLA: Compute-Efficient Pre-Training of LLMs via Low-Rank Activation

### v2

@@ Abstract
The full-size MLPs and the projection layers in attention introduce tremendous model sizes of large language models (LLMs), imposing extremely demanding needs of computational resources in the pre-training stage. However, we empirically observe that the activations of pre-trained LLMs exhibit low-rank property. Motivated by such observations, we propose CoLA and its memory-efficient implementation, CoLA-M, to replace these full-size layers with compute-efficient auto-encoders that naturally enforce low-rank activations throughout training. This fundamental architectural change eliminates the activation redundancy and significantly boosts model capacity and training efficiency. Experiments on LLaMA models with 60 million to 7 billion parameters show that CoLA reduces the computing cost by $\bf 2\boldsymbol{\times}$ and improves training throughput by $\bf 1.86\boldsymbol{\times}$ while maintaining full-rank level performance. CoLA-M further squeezes memory cost without sacrificing throughput, offering a pre-training approach with collectively superior parameter, computing, and memory efficiency. The LLMs produced are also $\bf 2\boldsymbol{\times}$ smaller, enabling faster inference with lower memory cost on resource-constrained platforms.  [^1] Code available here.
@@
Ziyue Liu*1, Ruijie Zhang*1, Zhengyang Wang*1, Zi Yang2, Paul Hovland3, Bogdan Nicolae3, Franck Cappello3, Zheng Zhang1 1University of California at Santa Barbara; 2University at Albany, SUNY 3Argonne National Laboratory
@@ 1 Introduction
Large foundation models have revolutionized the landscape of artificial intelligence, achieving unprecedented success in the language, vision, and scientific domains. In a quest to improve accuracy and capability, foundation models have become huge. Several studies Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024) have highlighted a rapid increase in the size of the model and the number of training tokens. Models such as 175B GPT-3 Brown et al. (2020), 405B LLaMA-3 Dubey et al. (2024), and 540B PaLM Chowdhery et al. (2023) are just a few examples of this trend. Under such circumstances, a large number of GPUs are needed in order to provide the computational and high-bandwidth memory capacity needed to pre-train large fundation models over long periods of time (months). The staggering increase in cost results in an unsustainable trend, prompting the need to develop cost-efficient pre-training techniques that reduce the scale, FLOPs, and GPU memory cost.
@@
Contributions: In this paper, we rethink the fundamental architecture of LLMs and propose CoLA: Compute-Efficient Pre-Training of LLMs via Low-rank Activation, and its memory efficient implementation CoLA-M, to achieve all the desirable properties mentioned above. We summarize our contributions as follows:
@@
Parameter $\boldsymbol{\downarrow}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Compute $\boldsymbol{\downarrow}$ Training $\boldsymbol{\times}$$\boldsymbol{\times}$
Inference $\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Memory $\boldsymbol{\downarrow}$ Training
Inference $\boldsymbol{\times}$$\boldsymbol{\times}$
Throughput $\boldsymbol{\uparrow}$ Training $\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Inference $\boldsymbol{\times}$$\boldsymbol{\times}$$\boldsymbol{\times}$
@@ 2 Related Work
Gradient Compression. GaLore Zhao et al. (2024) reduces memory by projecting gradients into a low-rank space, shrinking optimizer states below the typical 2×2$\times$2 × AdamW overhead Loshchilov (2017). However, it increases computation by adding up/down projections on top of already compute-heavy full-rank training. As shown in Fig. 1, its estimated FLOPs surpass full-rank training on the LLaMA-1B scale. Follow-up works Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024) further explore low-rank gradient projection. While being promising, these methods are mostly orthogonal to our focus. Crucially, they are computing lower-bounded by the full-rank baseline. Our goal instead is to reduce computing cost to a fraction of full-rank LLM pre-training.
Activation Compression. CompAct Shamshoum et al. (2024) reduces memory of the computational graph using low-rank compression on saved activations, which introduces similar computing cost, yet underperforms GaLore. ESPACE Sakr and Khailany (2024) explores a very similar idea by projecting activations based on well-trained weight matrices, thus only applicable to the post-training stage. Crucially, the projections in both methods introduce additional computing costs on top of the full-rank baseline. And both of them do not change the fundamental structure of LLMs.
This paper presents an architectural innovation that explicitly enforces low-rank activations by adopting the bottleneck-shaped auto-encoders as the building brick of the transformer architecture. This is conceptually different from the above model compression methods despite of some similarities in their formulations. Our approach is mostly orthogonal with gradient compression techniques, meaning that they could be combined to further boost efficiency.
@@ 3 CoLA for Efficient LLM Pre-Training
Figure 2: MLP Activation Spectrum of the pre-trained GPT-2 small Radford et al. (2019). Model activations are evaluated on the WikiText2 dataset. a) The singular value decay across different decoder blocks. b) The full dimension vs. effective rank ($\alpha=0.95$) per block.
Figure 3: Comparison between different efficient pre-training frameworks. a) LoRA/ReLoRA Lialin et al. (2023) freezes a full-rank weight; b) GaLore Zhao et al. (2024) only reduces optimizer states by down and up projecting gradients; c) SLTrain Han et al. (2024) requires reconstruction of the low-rank and sparse matrices; d) CoLA (ours) is a pure low-rank architecture involving only rank $r$ weight matrices.
@@ 3.1 A Motivating Example
Many works have observed the low-rank structure of model activations in deep neural networks Cui et al. (2020); Huh et al. (2021). We also observe this phenomenon in LLMs, i.e. the effective rank of the activations is much smaller than their original dimensionality. To quantify this, we define the effective rank $r(\alpha)$ of activation as the minimal number of singular values needed to preserve an $\alpha$-fraction of the total spectral energy. Formally:
$r(\alpha)\;=\;\min\left\{k\;\middle|\;\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{n}\sigma_{i}^{2}}\;\geq\;\alpha\right\},$ (1)
where $\sigma_{1},\sigma_{2},\ldots,\sigma_{n}$ are the singular values of the activation matrix, and $0<\alpha\leq 1$ is the desired ratio of preserved information. As shown in our experiments, the rapid decay of singular values [Fig. 2a] leads to much smaller $r(\alpha)$ compared to the full dimension [Fig. 2b]. This highlights the significant low-rank nature in the activations of pre-trained LLMs. More results showing the same pattern can be found in Appendix 11.
@@ 3.2 Low-Rank Activation via Auto-Encoder
The above observation motivates us to ask one fundamental question: do we really need these full-size MLP and linear layers in LLMs? To eliminate the redundant activations, we propose to replace them with bottleneck-structured auto-encoders that naturally facilitate low-rank activations.
Let $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ be the weight matrix of an arbitrary linear layer followed by a nonlinear activation in the transformer architecture:
$\mathbf{h}=\sigma\left(\mathbf{Wx}\right),\;\text{with}\;\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}.$ (2)
We replace this MLP layer with an auto-encoder layer which consists low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and a non-linear activation $\sigma$ in the middle. Rank $r<\min(d_{\text{in},\text{out}})$ is a design parameter that trades off between compute and performance. Formally, it can be written as:
$\mathbf{h^{\prime}}=\mathbf{B}\,\sigma(\mathbf{A}\mathbf{x}),$ (3)
@@
  • The auto-encoder is not equivalent to using smaller feature dimensions in MLP layers, since $\mathbf{B}$ in the current layer cannot be merged with $\mathbf{A}$ in the next layer, due to the existence of various operations (e.g. residual connection) in the original dimension.
Since the low-rank property is widely observed regardless of whether $\mathbf{Wx}$ being followed by nonlinearity (see details in Appendix 11), we also uniformly adopt this auto-encoder structure to all projection layers in the transformer architecture. We empirically find that adding the original nonlinearity on top of Eq. (3) does not harm the performance, nor necessarily brings benefit (c.f. Appendix 10).

### v3

@@ Abstract
The full-size MLPs and the projection layers in attention introduce tremendous model sizes of large language models (LLMs), consuming extensive computational resources in pre-training. We empirically observe that the activations of pre-trained LLMs exhibit low-rank property. Motivated by such observations, we propose CoLA and its memory-efficient implementation, CoLA-M, to replace these full-size layers with compute-efficient auto-encoders that naturally enforce low-rank activations throughout training. This fundamental architectural change eliminates the activation redundancy and significantly boosts model capacity and training efficiency. Experiments on LLaMA models with 60 million to 7 billion parameters show that CoLA reduces the computing cost by $\bf 2\boldsymbol{\times}$ and improves training throughput by $\bf 1.86\boldsymbol{\times}$ while maintaining full-rank level performance. CoLA-M further squeezes memory cost without sacrificing throughput, offering a pre-training approach with collectively superior parameter, computing, and memory efficiency. The LLMs produced are also $\bf 2\boldsymbol{\times}$ smaller, enabling faster inference with lower memory cost on resource-constrained platforms.  [^1] Code available here.
@@
Ziyue Liu*1, Ruijie Zhang*1, Zhengyang Wang*1, Mingsong Yan1, Zi Yang2, Paul Hovland3, Bogdan Nicolae3, Franck Cappello3, Sui Tang1, Zheng Zhang1 1University of California at Santa Barbara; 2University at Albany, SUNY 3Argonne National Laboratory {ziyueliu, ruijiezhang, zhengyangwang, zzhang01}@ucsb.edu
@@ 1 Introduction
Large foundation models have achieved unprecedented success in the language, vision, and scientific domains, but they have become huge. Several studies Kaplan et al. (2020); Hoffmann et al. (2022); Krajewski et al. (2024); Kumar et al. (2024) have highlighted a rapid increase in the size of the model and the number of training tokens. Models such as 175B GPT-3 Brown et al. (2020), 405B LLaMA-3 Dubey et al. (2024), and 540B PaLM Chowdhery et al. (2023) are just a few examples of this trend. Under such circumstances, a large number of GPUs are needed in order to provide the computational and high-bandwidth memory capacity needed to pre-train large fundation models over long periods of time (months). This unsustainable trend has prompted the need to develop cost-efficient pre-training techniques that reduce the scale, FLOPs, and GPU memory cost.
@@
Contributions: We rethink the fundamental architecture of LLMs and propose CoLA: Compute-Efficient Pre-Training of LLMs via Low-rank Activation, and its memory efficient implementation CoLA-M, to achieve all the desirable properties mentioned above. Our contributions include:
@@
  • We theoretically justify the benefit of using CoLA’s auto-encoder structure: they can be strictly better than conventional low-rank models under specific data-dependent conditions. We also derive an effective-rank–aware, non-asymptotic recovery bound that tightens as the spectrum concentrates.
@@
[…]
@@ 2 Related Work
Gradient Compression. GaLore Zhao et al. (2024) reduces memory by projecting gradients into a low-rank space, shrinking optimizer states below the typical $2\times$ AdamW overhead Loshchilov (2017). However, it adds up/down projections on top of already compute-heavy full-rank training. As shown in Fig. 1, its estimated FLOPs surpass full-rank training on the LLaMA-1B scale. Follow-up works Chen et al. (2024); Huang et al. ; Liao et al. (2024); Hao et al. (2024); Zhu et al. (2024) further explore low-rank gradient projection. While being promising, these methods are mostly orthogonal to our focus. Crucially, they are computing lower-bounded by the full-rank baseline. Our goal instead is to reduce computing cost to a fraction of full-rank LLM pre-training.
Activation Compression. CompAct Shamshoum et al. (2024) reduces memory of the computational graph using low-rank compression on saved activations, which slightly reduces the computing cost of Galore, yet underperforms both GaLore and full-rank training in terms of accuracy. ESPACE Sakr and Khailany (2024) explores a very similar idea by projecting activations based on well-trained weight matrices, thus only applicable to the post-training stage. Crucially, the projections in both methods introduce additional computing costs on top of the full-rank baseline. And both of them do not change the fundamental structure of LLMs.
This paper presents an architectural innovation that explicitly enforces low-rank activations by adopting the bottleneck-shaped auto-encoders as the building brick of the transformer architecture. This is conceptually different from the above model compression methods. Our approach is mostly orthogonal with gradient compression techniques, meaning that they could be combined to further boost efficiency.
@@ 3 CoLA for Efficient LLM Pre-Training
Figure 2: MLP activation [i.e., Eq.(2)] spectrum of the pre-trained GPT-2 small Radford et al. (2019). Model activations are evaluated on the WikiText2 dataset. a) The singular value decay across different decoder blocks. b) The full dimension vs. effective rank ($\alpha=0.95$).  [^3] We updated this figure to reflect the exact post-activation spectrum to avoid potential confusions in our original manuscript.
Figure 3: Comparison between different pre-training frameworks. a) LoRA/ReLoRA Lialin et al. (2023) freezes a full-rank weight; b) GaLore Zhao et al. (2024) only reduces optimizer states by down and up projecting gradients; c) SLTrain Han et al. (2024) requires reconstruction of the low-rank and sparse matrices; d) CoLA (ours) is a pure low-rank architecture involving only rank $r$ weight matrices.
@@ 3.1 A Motivating Example
Many works have observed the low-rank structure of model activations in deep neural networks Cui et al. (2020); Huh et al. (2021). We also observe this phenomenon in LLMs, i.e. the effective rank of the activations is much smaller than their original dimensionality. To quantify this, we define the effective rank of a matrix $\mathbf{C}$ as the minimal number of singular values needed to preserve an $\alpha$-fraction of the total spectral energy. Formally:
$r_{\alpha}(\mathbf{C})\;=\;\min\left\{k\;\middle|\;\frac{\sum_{i=1}^{k}s_{i}^{2}}{\sum_{i=1}^{n}s_{i}^{2}}\;\geq\;\alpha\right\},$ (1)
where $s_{1},s_{2},\ldots,s_{n}$ are the singular values of matrix $\mathbf{C}$, and $0<\alpha\leq 1$ is the desired ratio of preserved information. As shown in our experiments, the rapid decay of singular values [Fig. 3a] leads to much smaller effective ranks compared to the full dimension [Fig. 3b]. This highlights the significant low-rank nature in the activations of pre-trained LLMs. More results showing the same pattern can be found in Appendix 12.
@@ 3.2 Low-Rank Activation via Auto-Encoder
The above observation motivates us to ask one fundamental question: do we really need these full-size MLP and projection layers in LLMs? To eliminate the redundant activations, we propose to replace them with bottleneck-structured auto-encoders that naturally facilitate low-rank activations.
Let $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ be the weight matrix of an MLP layer: a linear layer followed by a nonlinear activation in the transformer architecture:
$\mathbf{h}_{\text{MLP}}=\sigma\left(\mathbf{Wx}\right),\;\text{with}\;\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}.$ (2)
We replace this MLP layer with an auto-encoder layer which consists of low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$ and $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and a non-linear activation $\sigma$ in the middle. Rank $r<\min(d_{\text{in},\text{out}})$ is a design parameter that trades off between compute and performance. Formally, it can be written as:
$\mathbf{h}_{\text{CoLA}}=\mathbf{B}\,\sigma(\mathbf{A}\mathbf{x}).$ (3)
We empirically find that adding the original nonlinearity on top of Eq. (3) does not harm or necessarily improve the accuracy (c.f. Appendix 11). Similarly, for linear layers that are not followed by an activation function, i.e., a projection layer in attention module (we continue using $\mathbf{W}$ for simplicity):
$\mathbf{h}_{\text{Linear}}=\mathbf{Wx},$ (4)
the low-rank property is also significantly present (see details in Appendix 12). Therefore, they are replaced by CoLA layers (3) as well.
@@
  • The auto-encoder is not equivalent to using smaller feature dimensions in MLP layers, since $\mathbf{B}$ in the current layer cannot be merged with $\mathbf{A}$ in the next layer, due to the existence of various operations (e.g. residual connection, element-wise product) in the original dimension.
@@
3.3 Theoretical Analysis
We theoretically justify the use of nonlinear activations in CoLA’s auto-encoders and offer an effective-rank–aware recovery bound. We first explain the benefit of CoLA over standard low-rank approximations on linear projection layers. Then we (partially) extend the analysis to MLP layers.
Let $n$ be the number of tokens, and $\sigma$ be a nonlinear activation function. Consider
$\mathcal{E}_{\sigma}(r):=\min_{\mathbf{A}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\mathbf{B}\in\mathbb{R}^{d_{\mathrm{out}}\times r}}\left\lVert\mathbf{Y-B\sigma(AX)}\right\rVert_{\mathrm{F}}.$ (5)
Here $\mathbf{X}\in\mathbb{R}^{d_{\mathrm{in}}\times n}$ denotes the input to the linear layer in the compressed network, or equivalently the output of the already–compressed preceding layers. The target $\mathbf{Y}\in\mathbb{R}^{d_{\mathrm{out}}\times n}$ denotes the original output at the same layer, typically $\mathbf{Y=WX_{\mathrm{True}}}$, where $\mathbf{X_{\mathrm{True}}}$ is the input to this layer of the originally uncompressed model. In general, $\mathbf{X}$ and $\mathbf{X_{\mathrm{True}}}$ need not coincide; they may differ through a (possibly nonlinear) transformation induced by the preceding layers and their compression. If the activation $\sigma$ is the identity, problem (5) reduces to the conventional low-rank method:
$\mathcal{E}_{\mathrm{id}}(r):=\min_{\mathbf{A}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\mathbf{B}\in\mathbb{R}^{d_{\mathrm{out}}\times r}}\mathbf{\left\lVert Y-BAX\right\rVert_{\mathrm{F}}}.$
The following proposition shows that the optimal value with the nonlinear activation is no larger than in the identity case. Full proofs are in Appendix G.
Proposition 3.1.
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, then $\mathcal{E}_{\sigma}(r)\leq\mathcal{E}_{\mathrm{id}}(r)$.
Let $\mathbf{\mathrm{row}(X)}$ denote the row space of $\mathbf{X}$. Under the identity activation, we notice that the approximation is confined to $\mathbf{\mathrm{row}(X)}$. The next result shows that, with a nonlinear activation, one can generate features $\mathbf{\sigma(u^{\top}X)}$ lying outside $\mathbf{\mathrm{row}(X)}$; hence, CoLA can represent outputs that are not realizable by standard low-rank approximation.
Proposition 3.2.
Suppose that $\mathbf{X}\in\mathbb{R}^{d_{\mathrm{in}}\times n}$ has no identical columns, no zero columns and satisfies $n>\mathrm{rank}(\mathbf{X})$. If $\sigma(0)=0$, $\sigma^{\prime}(0)\neq 0$ and $\sigma^{\prime\prime}(0)\neq 0$, then there exists $\mathbf{u}\in\mathbb{R}^{d_{\mathrm{in}}}$ such that $\sigma(\mathbf{u^{\top}X})$ is a nonzero vector and $\mathbf{\sigma(u^{\top}X)\notin\mathrm{row}(X)}$.
Next we identify a sufficient data-dependent condition under which the CoLA layer $\mathbf{B}\sigma(\mathbf{AX})$ strictly outperforms a standard low-rank layer $\mathbf{BAX}$, i.e., $\mathcal{E}_{\sigma}(r)<\mathcal{E}_{\mathrm{id}}(r)$. Informally, if rows of $\mathbf{Y}$ lie substantially outside $\mathbf{\mathrm{row}(X)}$ and align with a nonlinear feature $\mathbf{v^{\top}:=\sigma(u^{\top}X)\notin\mathrm{row}(X)}$, then CoLA will achieve a strictly better approximation to $\mathbf{Y}$. To ground our discussion, we introduce the following notations. Let $P_{\mathbf{X}}$ denote the orthogonal projector onto $\mathbf{\mathrm{row}(X)}$ and set $P_{\mathbf{X}^{\perp}}:=I-P_{\mathbf{X}}$, where $I$ is the identity operator. Define $\mathbf{Y}_{\parallel}:=P_{\mathbf{X}}\mathbf{Y}$ and $\mathbf{Y}_{\perp}:=P_{\mathbf{X}^{\perp}}\mathbf{Y}$ (projectors are applied row-wisely to matrices). Similarly, let $P_{\mathbf{v}}$ denote the orthogonal projector onto $\operatorname{span}\{\mathbf{v}^{\top}\}$, and define $P_{\mathbf{v}^{\perp}}:=I-P_{\mathbf{v}}$. For a matrix $\mathbf{Z}$ with singular values $s_{1}\geq s_{2}\geq\cdots$, where $s_{j}:=0$ for $j>\mathrm{rank}(\mathbf{Z})$, write $s_{>k}(\mathbf{Z}):=\bigl(\sum_{j>k}s_{j}^{2}\bigr)^{1/2}$.
Theorem 3.3.
Suppose that matrix $\mathbf{X}\in\mathbb{R}^{d_{\mathrm{in}}\times n}$ has no identical columns, no zero columns and satisfies $n>\mathrm{rank}(\mathbf{X})$. Suppose that $\sigma(0)=0$, $\sigma^{\prime}(0)\neq 0$ and $\sigma^{\prime\prime}(0)\neq 0$. Let $\mathbf{u}\in\mathbb{R}^{d_{\mathrm{in}}}$ and $\mathbf{v^{\top}:=\sigma(u^{\top}X)\notin\mathrm{row}(X)}$. If
$\left\lVert P_{\mathbf{v}^{\perp}}(\mathbf{Y})\right\rVert_{\mathrm{F}}^{2}<\left\lVert\mathbf{Y}_{\perp}\right\rVert_{\mathrm{F}}^{2}+\left(s_{>r}(\mathbf{Y}_{\parallel})\right)^{2},$ (6)
then $\mathcal{E}_{\sigma}(r)<\mathcal{E}_{\mathrm{id}}(r)$.
In the extreme case where each row of $\mathbf{Y}$ lies in $\mathrm{span}\{\mathbf{v}^{\top}\}$, assumption (6) holds trivially and $\mathcal{E}_{\sigma}(r)=0<\mathcal{E}_{\mathrm{id}}(r)$. We also note that (6) is sufficient (not necessary); sharper variants are possible.
We next assume that $\mathbf{Y}$ inherently admits an approximate autoencoder representation of $\mathbf{X}$, up to noise. The following theorem provides a non-asymptotic bound on the representation error of the $\mathcal{E}_{\sigma}(r)$ minimizer relative to this latent ground truth.
Theorem 3.4.
Suppose that there exist $\mathbf{A}_{\mathrm{True}}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\mathbf{B}_{\mathrm{True}}\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ such that
$\left\lVert\mathbf{Y}-\mathbf{B}_{\mathrm{True}}\sigma(\mathbf{A}_{\mathrm{True}}\mathbf{X})-\mathbf{G}\right\rVert_{\mathrm{2}}\leq\epsilon,$ (7)
where $\mathbf{G}$ is a random matrix of i.i.d Gaussian entries with zero-mean and variance $v^{2}$. Suppose that the optimal value $\mathcal{E}_{\sigma}(r)$ is obtained at $(\mathbf{A}^{*},\mathbf{B}^{*})$. Then with probability at least $1-2\exp(-(n+d_{\mathrm{out}}))$, it holds that
$\Delta:=\left\lVert\mathbf{B_{\mathrm{True}}\sigma(A_{\mathrm{True}}X)-B^{*}\sigma(A^{*}X)}\right\rVert_{\mathrm{F}}$
[…]
$\quad+s_{>r_{\alpha}(\mathbf{Y})}(\mathbf{Y})+\mathcal{E}_{\sigma}(r),$
where $\alpha\in(0,1]$, $s_{r_{\alpha}(\mathbf{Y})+1}$ is $(r_{\alpha}(\mathbf{Y})+1)$-th largest singular value of $\mathbf{Y}$ and $C$ is an absolute constant.
We note that the recovery bound explicitly depends on the effective rank $r_{\alpha}(\mathbf{Y})$, which is often empirically small (see Section 3.1). In particular, setting $\alpha=1$ reduces our result to a full-rank bound as $r_{1}(\mathbf{Y})=\mathrm{rank}(\mathbf{Y})$. When $\mathbf{Y}$ has a concentrated spectrum (i.e., $r_{\alpha}(\mathbf{Y})\ll\mathrm{rank}(\mathbf{Y})$), Theorem 3.4 generally yields a tighter bound than the full-rank case. In addition, the established error bound reflects the role of the nonlinear activation through the term $\mathcal{E}_{\sigma}(r)$. As shown in Theorem 3.3, under suitable conditions, $\mathcal{E}_{\sigma}(r)$ can be strictly smaller than its identity counterpart $\mathcal{E}_{\mathrm{id}}(r)$, thereby yielding a smaller overall error bound.
Partial Extension to MLP Layers. As stated in Section 3.2, we did not see a significant difference in performance when nonlinear activation was added on top of the auto-encoder layer [Eq (3)]. An auto-encoder followed by a non-linear activation is equivalent to just replacing the linear projection inside an MLP layer with Eq (3), therefore our above theoretical analysis still holds. We still need more theoretical understanding of the case without activation after Eq (3), which will be a future work.